Jose can paint a room in two hours, juanita can paint the room in three hours, how long would it take them to paint the room working together?
Answer by nerdybill(7384)

Jose can paint a room in two hours, juanita can paint the room in three hours, how long would it take them to paint the room working together?
.
Let x = time (hours)it takes for both
then
x(1/2 + 1/3) = 1
multiplying both sides by 6:
x(3 + 2) = 6
x(5) = 6
x = 6/5 hours
or
x = 1 hour and 12 minutes